### v1

@@ 5 Numerical experiments
We compare the performance of IRKGL16-SIMD with several state-of-the-art explicit symplectic integrators for the high-precision integration of non-stiff Hamiltonian systems using constant step size. A range of representative examples is considered to assess both accuracy and efficiency.
Some of the explicit symplectic integrators used in the comparison are available through the Julia language’s DifferentialEquations.jl suite [16]. However, for consistency and fair benchmarking, we have implemented all the reference integrators described in the previous section. We verified that, for the problems considered, our implementations of the splitting methods are at least as efficient as—and often more so than—those available in DifferentialEquations.jl.
The source code for the integrators, along with all Jupyter notebooks used to generate the numerical results, is publicly available at: https://github.com/mikelehu/Implicit_symplectic_can_outperform_explicit_symplectic.
All numerical experiments were conducted on an 11th Gen Intel Core i7-11850H processor (2.5 GHz, 16 threads), a CPU equipped with 512-bit SIMD registers. We used the long-term support (LTS) stable release of Julia version 1.10.7 (November 26, 2024) for all tests. To optimize SIMD performance, Julia was launched from the terminal with bounds checking disabled using the flag --check-bounds=no.
@@ 5.1 Test problems
We have made comparisons with systems of first-order differential equations as well as with systems of second-order differential equations.
@@ 5.1.1 System of first-order differential equations
Although [15] provides analytical expressions for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, their direct implementation is computationally inefficient. To enable a fair and representative comparison with our implicit symplectic integrator (IRKGL16), we have improved the evaluation of these flows, significantly enhancing the performance of the explicit method. The derivation and implementation details of this optimized explicit integrator are presented in Appendix A.
As in [15], we set the parameters to $E=0.995$, $L=4.6$, and $\beta=8.9\times 10^{-4}$. The initial conditions are $\theta=\pi/2$ and $p_{r}=0$, with $r=11$ chosen so that the corresponding orbit is regular (i.e., lies on a closed curve). The initial value of $p_{\theta}>0$ is determined by imposing the energy condition $H=-\tfrac{1}{2}$.
@@ 5.1.2 Systems of second-order differential equations
We examine a simplified model of the outer Solar System, which includes the Sun, the four outer planets (Jupiter, Saturn, Uranus, and Neptune), and Pluto, interacting via mutual Newtonian gravity. The system is described by an 18-degree-of-freedom Hamiltonian, with $q_{i},p_{i}\in\mathbb{R}^{3}$ for $i=1,\dots,6$, and Hamiltonian function
@@
5.2 Results of Numerical Integrations
5.2.1 Systems of First-Order Differential Equations
Figure 1 presents work-precision diagrams for the Schwarzschild black hole problem. The left panel shows the maximum local error in the Hamiltonian,
$\Delta H^{\text{loc}}_{\max}=\max_{n}\left(\Delta H^{\text{loc}}_{n}\right),\quad\Delta H^{\text{loc}}_{n}=\left|\frac{H(y_{n})-H(y_{n-1})}{H(y_{n-1})}\right|,$ (28)
as a function of the total number of function evaluations. The right panel plots the same error against CPU time. In this plot, the minimal CPU time for the IRKGL16-SIMD method is highlighted with a red vertical dashed line.
We observe that even the sequential implementation of IRKGL16 outperforms the explicit symplectic integrators for the high precision integration of this problem. As for the SIMD implementation, for the same CPU time budget, it achieves significantly higher accuracy even for lower precision integrations.
@@ 5.2.1 Systems of First-Order Differential Equations
Figure 1: Work-precision diagrams (maximum local error in the Hamiltonian) for the Schwarzschild black hole problem using the IRKGL16 method and symmetric composition schemes of orders $r=2,4,6,8$, and $10$. In the right plot, the sequential IRKGL16 implementation is denoted as IRKGL16-SEQ
Figure 2 shows the evolution of error over time for different integrators. Each method is integrated using a step size chosen such that the corresponding CPU time approximately matches the red vertical dashed line in the right panel of Figure 1. The left panel displays the global error in the Hamiltonian,
@@
while the right panel shows the relative error in the radial coordinate:
$\Delta r_{n}=\left|\frac{r_{n}-\tilde{r}_{n}}{\tilde{r}_{n}}\right|,$
where $\tilde{r}_{n}$ denotes the reference solution obtained using IRKGL16-SIMD with a comparatively small step size ($h=16$).
Figure 2: Left: global error in the Hamiltonian. Right: relative error in the radial coordinate $r$ for the Schwarzschild black hole problem. Results are shown for IRKGL16-SIMD and symmetric composition methods of orders $r=2,4,6,8,10$.
5.2.2 Systems of Second-Order Differential Equations
Figure 3 shows work-precision diagrams for the two second-order Hamiltonian problems. The top row corresponds to the 6-body outer Solar System model, while the bottom row pertains to the Hénon-Heiles system. The left panels display the maximum local Hamiltonian error (28) versus function evaluations, and the right panels show the same error versus CPU time. The red dashed line marks the optimal CPU time for the IRKGL16-SIMD method, in the sense that it is the minimal time required to achieve a local accuracy close to round-off error levels.
Figure 3: Work-precision diagrams (maximum local Hamiltonian error) for the 6-body outer Solar System and the Hénon-Heiles problem using IRKGL16, symmetric compositions (orders $r=2,4,6,8,10$), and symplectic RKN methods of orders 6 and 8.
Figures 46 show the evolution of errors, using simulations performed at optimal CPU time. The left panel in Figure 4 corresponds to the global Hamiltonian error of the 6-body outer Solar System, and the right panel to the Hénon-Heiles problem.
Figure 4: Time evolution of the global Hamiltonian error for the 6-body outer Solar System (left) and the Hénon-Heiles problem (right), using IRKGL16, symmetric compositions, and RKN methods.
Figure 5 depicts the position error of each planet in the outer Solar System over time, comparing IRKGL16 and SS05 ($r=10$). Figure 6 shows the error in both position and momentum for the Hénon-Heiles system across the different integrators.
Figure 5: Evolution of position error for each planet in the 6-body outer Solar System using IRKGL16 and SS05 ($r=10$).
@@ 5.2.2 Systems of Second-Order Differential Equations
Figure 6: Evolution of error in position (left) and momentum (right) for the Hénon-Heiles problem using IRKGL16, symmetric compositions, and RKN methods.

### v2

@@ 5 Numerical experiments
We compare the performance of IRKGL16 with several state-of-the-art explicit symplectic integrators for the high-precision integration of non-stiff Hamiltonian systems using constant step size. A range of representative examples is considered to assess both accuracy and efficiency.
It is well known [13] that, for long-term integrations, it is advisable to apply some form of improved summation (e.g., Kahan’s compensated summation [15]) in (17) to prevent excessive accumulation of roundoff errors. The idea is to increase the accuracy of the sum by representing each component of $y_{n}$ (for $n\geqslant 0$) as the sum of two floating-point numbers while representing each component of the update $\sum_{i=1}^{n}L_{n,i}$ (expected to be of relatively smaller size) as a floating-point number. In IRKGL16, we employ Julia’s Base.TwicePrecision for this purpose, which provides greater accuracy (though at higher computational cost) than Kahan’s compensated summation.
Some of the explicit symplectic integrators used in our comparison are available through the Julia DifferentialEquations.jl suite [19], but no compensated summation is applied there. For consistency and fair benchmarking, we implemented in Julia all the reference integrators described in the previous section using (optionally) Kahan’s compensated summation. We verified that, for the problems considered, our implementations of the splitting methods are at least as efficient as (and often more efficient than) the implementations in DifferentialEquations.jl when both are used without compensated summation.
The source code for the integrators, as well as all Jupyter notebooks used to reproduce the numerical results, is publicly available in the repository associated with this article [3].
All numerical experiments were conducted on an 11th Gen Intel Core i7-11850H processor (2.5 GHz, 16 threads) equipped with 512-bit SIMD registers. We used the long-term support (LTS) stable release of Julia v1.10.10 (June 27, 2025) for all tests.
@@ 5.1 Test problems
We have made comparisons for a Hamiltonian system of first-order differential equations split into three solvable parts, and two Hamiltonian systems of second-order differential equations with Hamiltonian (8) separable in two solvable parts.
@@ 5.1.1 System of first-order differential equations
Although [18] provides analytical expressions for the $t$-flow maps $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, their direct implementation is computationally inefficient. To enable a fair and representative comparison with our implicit symplectic integrator (IRKGL16), we have improved the evaluation of these $t$-flows, significantly enhancing the performance of the explicit method. Unlike the analytical expressions provided in [18] for the $t$-flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, our implementation avoids the use of trigonometric and inverse trigonometric functions entirely, resulting in improved computational efficiency. The derivation and implementation of the optimized flows are provided in the documentation available in the software repository [3] associated with this article.
As in [18], we set the parameters to $E=0.995$, $L=4.6$, and $\beta=8.9\times 10^{-4}$. The initial conditions are $\theta=\pi/2$ and $p_{r}=0$, with $r=11$ chosen so that the corresponding orbit is regular (i.e., lies on a closed curve). The initial value of $p_{\theta}>0$ is determined by imposing the energy condition $H=-\tfrac{1}{2}$. The integration is performed over the interval $0,10^{5}]$.
@@ 5.1.2 Systems of second-order differential equations
We examine a simplified model of the outer Solar System, which includes the Sun, the four outer planets (Jupiter, Saturn, Uranus, and Neptune), and Pluto, as point masses interacting via mutual Newtonian gravity. The system is described by an 18-degree-of-freedom Hamiltonian, with $q_{i},p_{i}\in\mathbb{R}^{3}$ for $i=1,\dots,6$, and Hamiltonian function
@@
5.2 Efficiency comparison of different implementations of $s$-stage IRKGL schemes
In Figure 1 we show work–precision diagrams for several implementations of $s$-stage IRKGL schemes applied to the Schwarzschild black hole problem.
Figure 1: (a) Maximum local Hamiltonian error versus total function evaluations for the $s$-stage IRKGL schemes ($s=2,4,6,8$: IRKGL4, IRKGL8, IRKGL12, IRKGL16). (b) The same error versus CPU time for two implementations of each scheme: fully vectorized (simd=true, fseq=false) and fully sequential (simd=false, fseq=true)
The left panel of each diagram reports the maximum local error in the Hamiltonian,
$\Delta H^{\mathrm{loc}}_{\max}=\max_{n}\left(\Delta H^{\mathrm{loc}}_{n}\right),\qquad\Delta H^{\mathrm{loc}}_{n}=\left|\frac{H(y_{n})-H(y_{n-1})}{H(y_{n-1})}\right|,$ (39)
as a function of the total number of function evaluations, for our implementations of the $s$-stage IRKGL schemes with $s=2,4,6,8$ (IRKGL4, IRKGL8, IRKGL12, and IRKGL16, respectively). This provides a measure of the intrinsic relative efficiency of the methods, independent of the computational environment, under the assumptions that (i) the user-supplied function that evaluates the right-hand side of the ODE system is executed once per stage vector, in sequential order, and (ii) the total CPU time is dominated by these function evaluations. According to this metric, IRKGL16 is the most efficient method in the high-accuracy regime, whereas IRKGL12 performs slightly better than IRKGL16 at moderate accuracy levels.
The right panel plots the same error against actual CPU time for two versions of each $s$-stage IRKGL method: the fully vectorized implementation (simd=true, fseq=false) and the fully sequential one (simd=false, fseq=true). The relative performance of the fully sequential versions is consistent with the behavior observed in the left panels. In contrast, the fully vectorized implementation of IRKGL16 outperforms the lower-order schemes even at moderate accuracy. This improvement is due to the higher SIMD speed-up achieved by the 8-stage scheme.
Figure 2: Panels (a)–(b) correspond to the 6-body outer Solar System, and panels (c)–(d) to the Hénon–Heiles system. (a, c) Maximum local Hamiltonian error versus total function evaluations for the $s$-stage IRKGL schemes ($s=2,4,6,8$). (b, d) The same error versus CPU time for two implementations: fully vectorized and fully sequential
In Figure 2, we present analogous work–precision diagrams for the two second-order Hamiltonian problems. The results again indicate that the fully vectorized implementation of IRKGL16 is the most efficient integrator for both moderate and high accuracy levels.
5.3 Numerical comparisons of IRKGL16 with explicit symplectic integrators
5.3.1 Work–precision diagrams
Figure 3: Work-precision diagrams (maximum local Hamiltonian error versus number of evaluations of $g(t,u)$) for the 6-body outer Solar System and the Hénon–Heiles problem, comparing IRKGL16 with symmetric composition schemes (orders $r=2,4,6,8,10$) and symplectic RKN methods of orders 6 and 8
We now present work–precision diagrams for the two second-order Hamiltonian problems considered in Subsection 5.1.2. The computational cost of one step of any splitting method (35) applied with the flows (32) is dominated by the $s$ applications of the $t$-flow $\varphi^{B}_{t}(q,v)\mapsto(q,v+t\,g(q))$, where $g(q)=M^{-1}\nabla U(q)$. Thus, for splitting methods, the relative cost of a numerical integration can essentially be measured by the number of calls to the function $g(u)$. For implicit methods applied to Hamiltonian second-order ODEs of the form (9), the relative cost can also be quantified in terms of function evaluations of the right-hand side $g(q)$, provided that it is sufficiently expensive to compute.
Figure 3 presents work–precision diagrams for the two second-order ODE problems, displaying the maximum local error in the Hamiltonian (39) plotted against the number of function evaluations for several splitting methods and IRKGL16. We consider the splitting integrators obtained by applying the composition schemes of the previous subsection with (31) as a basic integrator, and additionally the two RKN-type splitting methods of orders 6 and 8 selected in Section 4, all implemented with compensated summation.
From the efficiency diagrams shown in Figure 3, the larger overhead inherent to implicit schemes implies that the best explicit symplectic methods will outperform sequential implementations of IRKGL16 at low and moderate accuracy levels, regardless of the programming language or other aspects of the computational environment. However, at higher target accuracies, i.e., when the maximum local energy error is just above roundoff level, a sequential implementation of IRKGL16 may surpass explicit methods, provided that its overhead remains sufficiently small. This situation frequently arises in implementations of interpreted programming languages such as MATLAB: although the user-defined evaluation of the ODE right-hand side is performed in MATLAB, the linear-algebra operations responsible for most of the overhead are executed by highly optimized underlying routines [19, 12].
However, in other programming environments such as the Julia language, the work–precision diagrams based on function evaluations may not reflect the true computational efficiency.To quantify this, we introduce the CPU time per function evaluation ratio (CPU-FCN ratio for short), defined as the ratio of the total CPU time required for a given integration to the total number of function evaluations performed during that integration. For explicit splitting methods, this ratio is essentially constant: it depends neither on the step size nor on the particular scheme (i.e., the number of flows evaluated per step), since the computational cost is entirely dominated by the function evaluations themselves. For IRKGL16, by contrast, each step involves three additional sources of overhead beyond the function evaluations: the initial extrapolation to provide a starting guess for the fixed-point iteration, the overhead of the fixed-point iterations themselves, and the updating of state variables at the end of the step. Table 1 reports, for each of the second-order ODE problems, the CPU-FCN ratio of IRKGL16 normalized by that of the explicit splitting methods, for both the fully sequential implementation IRKGL16-SEQ (simd=false, fseq=true) and the fully vectorized implementation IRKGL16-SIMD (simd=true, fseq=false). For IRKGL16-SEQ, the normalized ratio exceeds two in both test problems, indicating that the cumulative overhead surpasses the cost of the function evaluations themselves. For IRKGL16-SIMD, by contrast, the normalized ratio is well below one. This is because the SIMD vectorization not only reduces the overhead of the implicit solver, but also enables the 8 stage function evaluations within each fixed-point iteration to be executed in parallel across SIMD registers, effectively amortizing their cost. As a consequence, each function evaluation in IRKGL16-SIMD is significantly cheaper in terms of wall-clock time than in the explicit methods, where function evaluations are performed sequentially. This implies that, relative to the work–precision diagrams presented above, the true efficiency curve of IRKGL16-SIMD is shifted to the left, making it even more competitive against the explicit symplectic integrators than those diagrams suggest.
Table 1: CPU-FCN ratio of IRKGL16 normalized by that of the explicit splitting methods, for the two second-order ODE test problems. “6-body Solar System (vec)” denotes the SIMD-vectorized implementation of the flows
IRKGL16-SEQ IRKGL16-SIMD
Hénon–Heiles 2.5 0.6
6-body Solar System 2.6 0.6
6-body Solar System (vec) 3.5 0.8
It should be noted that the CPU-FCN ratio of IRKGL16 is not constant across step sizes. While the overhead of the fixed-point iterations scales proportionally with the number of function evaluations per step, the overhead of the initial extrapolation and the state update does not. Consequently, for smaller step sizes—which require fewer fixed-point iterations per step—these fixed costs carry more relative weight, leading to a slightly larger CPU-FCN ratio. The values reported in Table 1 are computed at a reference step size for which the relative local energy error is approximately $10^{-14}$.
It is worth noting that explicit splitting methods can also benefit from SIMD vectorization, by explicitly exploiting SIMD registers in the implementation of the individual flows. To illustrate this, we have implemented a SIMD-optimized version of the flows for the 6-body outer Solar System problem, which results in a reduced CPU-FCN ratio for the explicit methods. We remark that SIMD-optimizing the flows requires additional implementation effort from the user, whereas the SIMD vectorization in IRKGL16 is handled internally with a generic implementation of the user-defined function $g$.
To assess the impact of this optimization on the relative computational cost of IRKGL16, Table 1 also reports (last row) the CPU-FCN ratio of IRKGL16 normalized by that of the SIMD-optimized explicit implementation. As expected, the normalized ratios increase with respect to the non-vectorized case, reflecting the fact that the explicit methods now perform each function evaluation more cheaply. Nevertheless, IRKGL16-SIMD retains a normalized ratio well below one, confirming that its fully vectorized implementation remains competitive even when compared against SIMD-optimized explicit integrators.
@@ 5.3.1 Work–precision diagrams
Figure 4: Work-precision diagrams (maximum local Hamiltonian error versus CPU time) for the Schwarzschild black hole problem using symmetric composition schemes of orders $r=2,4,6,8$, and $10$, and two variants of IRKGL16: (i) fully vectorized (IRKGL16-SIMD), (ii) fully sequential (IRKGL16-SEQ)
Figure 4 presents work–precision diagrams for the Schwarzschild black hole problem, displaying the maximum local error in the Hamiltonian (39) plotted against CPU time for explicit symplectic methods of orders $2,4,6,8,10$ and for two variants of IRKGL16: (i) the fully vectorized implementation IRKGL16-SIMD (simd=true, fseq=false), and (ii) the fully sequential implementation IRKGL16-SEQ (simd=false, fseq=true). In this example, we plot directly against CPU time rather than function evaluations, since the evaluation of the $t$-flows required to compute the basic map $\phi_{h}$ in (33) is considerably more expensive than the evaluation of the right-hand side of the ODE system, making the number of function evaluations a poor proxy for computational cost.
We observe that IRKGL16 outperforms the explicit symplectic integrators across the entire accuracy range shown, with IRKGL16-SEQ already surpassing them in the high-accuracy regime. A key factor is that the $t$-flow evaluations in (33) are considerably more expensive than the right-hand side evaluations in IRKGL16. As a result, for any given CPU-time budget, IRKGL16-SIMD delivers substantially higher accuracy than the explicit methods, regardless of the accuracy level they achieve.
5.3.2 Evolution of errors for the high-precision regime
Figure 5: Left: global error in the Hamiltonian. Right: relative error in the radial coordinate $r$ for the Schwarzschild black hole problem. Results are shown for IRKGL16-SIMD and symmetric composition methods of orders $r=2,4,6,8,10$
We now examine the time evolution of both the global Hamiltonian error and the position error for each of the three test problems in the high-precision regime. For each problem, we select a reference step size $h_{\mathrm{ref}}$ for IRKGL16-SIMD such that the maximum global Hamiltonian error
@@ 5.3.2 Evolution of errors for the high-precision regime
is approximately $10^{-14}$. Each explicit method is then run with a step size chosen so that its total CPU time matches that of IRKGL16-SIMD at $h_{\mathrm{ref}}$. This ensures a fair comparison on equal computational budget grounds.
For the Schwarzschild black hole problem, the results are shown in Figure 5. The left panel displays the global Hamiltonian error and the right panel the relative error in the radial coordinate $r$, defined as $\Delta r_{n}=|r_{n}-\tilde{r}_{n}|/|\tilde{r}_{n}|$, where $\tilde{r}_{n}$ denotes the reference solution computed with IRKGL16-SIMD at a smaller step size ($h=h_{\mathrm{ref}}/4$). We compare IRKGL16-SIMD against symmetric composition methods of orders $r=2,4,6,8,10$. At the CPU-time-matched step sizes, the most competitive explicit methods are those of orders 2 and 4, yet their global Hamiltonian and position errors remain approximately nine orders of magnitude larger than those of IRKGL16-SIMD throughout the integration.
Figure 6 presents the analogous comparison for the Hénon-Heiles problem, showing the global Hamiltonian error (left) and the absolute error in position $q$ (right), with the reference solution again provided by IRKGL16-SIMD at small step size. Here the most competitive explicit methods at equal CPU cost are the RKN-type integrators BM02 (order 6) and BCE22 (order 8), yet IRKGL16-SIMD maintains a substantially lower error in both the Hamiltonian and position throughout the integration interval.
For the 6-body outer Solar System, the comparison is carried out against SIMD-optimized implementations of the two most competitive 8th-order explicit integrators, SS05r8 and BCE22, with step sizes matched to the CPU time of IRKGL16-SIMD using the SIMD-optimized evaluation of the acceleration function $g$. This represents a particularly challenging benchmark for IRKGL16, as the explicit methods benefit from problem-specific SIMD acceleration. Figure 7 shows the time evolution of the global Hamiltonian error: even under these conditions, IRKGL16-SIMD achieves lower Hamiltonian errors than both explicit competitors throughout the integration. The evolution of position errors for each planet is shown in Figure 8, comparing IRKGL16-SIMD (left) against SS05r8 (right), which exhibits lower position errors than BCE22; again, IRKGL16-SIMD yields lower position errors across all planets over the entire integration interval.
@@
Figure 6: Hénon-Heiles problem: time evolution of (i) the global Hamiltonian error (left), (ii) error in position (right) using IRKGL16, symmetric compositions, and RKN methods
Figure 7: Time evolution of the global Hamiltonian error for the outer Solar System 6-body problem, computed with IRKGL16 and the 8th-order integrators SS05r8 and BCE22 using SIMD-vectorization of the flows.
Figure 8: Evolution of the position error for each planet in the outer Solar System 6-body problem: IRKGL16 (left) and the 8th-order integrator SS05r8 using SIMD-vectorization of the flows (right).
Across all three test problems considered—the Schwarzschild black hole, the Hénon-Heiles system, and the 6-body outer Solar System—IRKGL16-SIMD consistently achieves lower global Hamiltonian and position errors than the best-performing explicit symplectic integrators at equal CPU time in the high-precision regime.
5.4 Discussion on alternative implementations
The main source of overhead in implicit methods implemented with fixed-point iterations arises from the repeated updating of the state variables stored in the $D$-dimensional arrays $U_{i}$ (and $L_{i}$) at each iteration. At least two distinct approaches can be adopted to exploit SIMD registers in order to reduce this overhead.
The first approach performs the updates of each state variable in the $D$-dimensional arrays $U_{i}$ independently, and vectorizes them component-wise, provided that the state components of each array $U_{i}$ are stored consecutively in memory. This is the more standard approach, and can in principle be handled automatically by the compiler, provided that the code is written appropriately. Specifically, the inner loops over $j=1,\dots,s$ required for the stage summations (16) or (20) must be unrolled, so that the outer loop over $k=1,\dots,D$ consists of straightforward arithmetic that the compiler can SIMD-vectorize. We have verified that, in our sequential implementation of IRKGL16, unrolling these inner loops and disabling bounds checking via the @inbounds macro causes the Julia compiler to apply component-wise vectorization automatically, substantially reducing the overhead. A further, marginal reduction can be achieved by applying, instead of @inbounds, the @turbo macro from the LoopVectorization.jl package.
The second approach instead arranges, for each state variable, the corresponding components $U_{1},\dots,U_{s}$ consecutively in memory, and vectorizes their simultaneous update using SIMD $s$-vectors. This is the approach adopted in our implementation of IRKGL16, explained in detail in Subsection 2.5 and Section 3. We have verified that when the right-hand side function evaluations are performed sequentially, both vectorization approaches—the stage-wise vectorization of IRKGL16 with simd=true, fseq=false on one hand, and the component-wise vectorization via unrolled inner loops with the @turbo macro on the other—yield essentially the same reduction in overhead. The key advantage of stage-wise vectorization is that it enables the $s$ evaluations of the right-hand side function to be performed in parallel within SIMD registers, whenever it is implemented as a generic function compatible with the SIMD.jl package. This SIMD-parallel evaluation is handled internally by IRKGL16 and remains completely transparent to the user, who is only required to provide a standard generic implementation of the right-hand side function. In summary, we advocate for stage-wise vectorization —and have adopted it in IRKGL16 in particular—as a convenient and user-friendly implementation strategy for IRKGL schemes.
To further assess the practical relevance of component-wise vectorization, we computed the CPU-FCN ratios of the Fortran 77 implementations of both the 12th-order IRKGL scheme and the explicit composition schemes provided in the gnicodes folder of the software repository accompanying [13], applied to the 6-body outer Solar System problem and compiled with the -O2 and -O3 optimization flags, the latter requesting vectorization whenever possible. We do not consider the -Ofast flag, since it enables aggressive optimization that silently eliminates compensated summation algorithms by treating their error-compensation steps as algebraically redundant, thereby compromising the long-term energy error propagation control that symplectic integrators rely on. For the explicit composition schemes, the CPU-FCN ratio normalized by that of our SIMD-optimized Julia implementation of the flows is approximately 1.0 under -O3 and 1.2 under -O2, suggesting that the compiler successfully applies SIMD acceleration under -O3. This also shows that similar efficiency can be achieved for explicit symplectic schemes in both Fortran and Julia, provided that the implementation is set up so that the compiler can exploit SIMD acceleration. For IRKGL12, by contrast, the normalized CPU-FCN ratio is approximately 1.4 under both -O2 and -O3, indicating that the compiler fails to apply vectorization even under -O3. The reason is that the number of stages $s$ is treated as a generic parameter in the gnicodes implementation, which prevents the inner summation loops from being unrolled. We have verified that setting $s=6$ statically, so that the number of terms in the summation is known at compile time, allows the compiler to unroll the loop and successfully apply component-wise vectorization under -O3, reducing the normalized CPU-FCN ratio to approximately 1.05 (compared to 1.4 under -O2). This confirms that automatic component-wise vectorization is also achievable in Fortran implementations of IRKGL schemes, provided the number of stages is fixed at compile time. However, even in this favorable case, the normalized CPU-FCN ratio remains above one, since no parallelism is exploited across the $s$ independent evaluations of the right-hand side function. By contrast, IRKGL16-SIMD achieves a normalized CPU-FCN ratio of approximately 0.8 in the 6-body Solar System problem (normalized against explicit schemes with SIMD-optimized right-hand-side function), by seamlessly and transparently exploiting stage-wise SIMD parallelism across all function evaluations.
For ODE systems whose right-hand side is not compatible with SIMD.jl, an alternative strategy is to evaluate the $s$ stage functions in parallel using multithreading. In Julia, this can be conveniently implemented using the Threads.@threads macro. However, the additional synchronization overhead incurred by thread-level parallelism is only compensated when the right-hand side function is sufficiently expensive to evaluate [14]. We have verified that, for the three test problems considered in this work, a straightforward application of Threads.@threads to parallelize the stage function evaluations results in increased total CPU time, indicating that the right-hand side functions in these examples are not costly enough to amortize the threading overhead.